OpenStudy (anonymous):

need help asap questions are in reply area and testimonial fan and medl for whoever helps

OpenStudy (anonymous):

1. Factor x2 + 6x + 8. (Points : 1) (x + 1)(x + 6) (x + 1)(x + 8) (x + 2)(x + 4) prime Question 2. 2. Factor x2 + 7x + 12. (Points : 1) (x + 3)(x + 4) (x + 1)(x + 7) (x + 2)(x + 6) prime Question 3. 3. Factor x2 + 10x – 18. (Points : 1) (x – 2)(x – 9) (x + 3)(x – 6) (x – 2)(x + 5) prime Question 4. 4. Factor x2 – 10x + 24. (Points : 1) (x – 2)(x – 12) (x – 3)(x – 8) (x – 4)(x – 6) prime Question 5. 5. Factor x2 – 2x – 24. (Points : 1) (x – 6)(x + 4) (x + 6)(x – 4) (x – 12)(x + 2) (x + 12)(x – 2)

OpenStudy (usukidoll):

@dan815 yo man help!!!!

OpenStudy (dan815):

https://www.youtube.com/watch?v=ZQ-NRsWhOGI HOW TO FACTOR

OpenStudy (usukidoll):

=/

OpenStudy (dan815):

(x + 2)(x + 4) , expand this what do u get

OpenStudy (dan815):

(x + 2)(x + 4) = (x + 2)(x) + (x + 2)(4) = x*x + 2*x + x*4+2*4 =?

OpenStudy (dan815):

simplify more

OpenStudy (usukidoll):

I prefer FOIL xD

OpenStudy (usukidoll):

First outer inner last